# Towards Anomaly-Aware Pre-Training and Fine-Tuning for Graph Anomaly Detection

### v2

@@ 3.3 Theoretical Insights
In this subsection, we provide a theoretical analysis to support our architectural design. Our theoretical analysis is grounded in the Contextual Stochastic Block Model (CSBM) CSBM , a widely used generative model for attributed graphs GCTheory ; HomoNece ; Demy ; Node-MoE ; HSBM . To properly reflect the homophily disparity, we introduce a novel variant, the Anomalous Stochastic Block Model (ASBM), which combines homophilic and heterophilic structures within a unified generative framework:
Definition 1 ($ASBM(n_{a},n_{n},\boldsymbol{\mu},\boldsymbol{\nu},(p_{1},q_{1}),(p_{2},q_{2}),P_{a},P_{n})$).
ASBM generates two disjoint node sets: abnormal nodes $\mathcal{C}_{a}$ and normal nodes $\mathcal{C}_{n}$, with sizes $n_{a}$ and $n_{n}$, respectively. Feature matrices $\boldsymbol{X}_{a}\in\mathbb{R}^{n_{a}\times d}$ and $\boldsymbol{X}_{n}\in\mathbb{R}^{n_{n}\times d}$ are sampled row-wise from $\mathcal{N}(\boldsymbol{\mu},\frac{1}{d}\boldsymbol{I})$ and $\mathcal{N}(\boldsymbol{\nu},\frac{1}{d}\boldsymbol{I})$, where $|\boldsymbol{\mu}|_{2},|\boldsymbol{\nu}|_{2}\leq 1$. The full feature matrix $\boldsymbol{X}$ is formed by stacking $[\boldsymbol{X}_{a};\boldsymbol{X}_{n}]$. Each node can follow either a homophilic or heterophilic connectivity pattern. In the homophilic regime, edges are generated with intra-class probability $p_{1}$ and inter-class probability $q_{1}$, where $p_{1}>q_{1}$; in the heterophilic regime, $p_{2}<q_{2}$. We further denote $\mathcal{H}_{o}$ and $\mathcal{H}_{e}$ as the set of homophilic and heterophilic nodes and assume that all nodes share a common expected degree: $p_{1}+q_{1}=p_{2}+q_{2}$.
To examine the impact of different spectral filters on ASBM, following prior analyses GCTheory ; Node-MoE , we adopt a linear classifier parameterized by $\boldsymbol{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$. Here, we denote $\boldsymbol{\hat{X}}=\boldsymbol{U}g(\boldsymbol{\Lambda})\boldsymbol{U}^{\top}\boldsymbol{X}$ as the frozen filtered features, serving as fixed representations passed to downstream classifiers. Then, the predicted labels are given by $\boldsymbol{\hat{y}}=\sigma(\boldsymbol{\hat{X}}\boldsymbol{w}+b\boldsymbol{1})$, where $\sigma(\cdot)$ is sigmoid function. This classifier is optimized using the trivial binary cross-entropy loss defined in Eq. (12). The separability of this model under node-adaptive filtering is characterized by the following theorem:
@@
Assume the graph size $n=n_{a}+n_{n}$ is relatively large with $\omega(d\log d)\leq n\leq\mathcal{O}(poly(d))$, the graph is not too sparse with $p_{1},q_{1},p_{2},q_{2}=\omega(\log^{2}(n)/n)$, the feature center distance is not too small with $\|\boldsymbol{\mu}-\boldsymbol{\nu}\|=\Omega(\frac{\log n}{\sqrt{dn(p_{1}+q_{1})}})$, the sizes of the homophilic and heterophilic node sets are approximately equal, and the norm of $\|\boldsymbol{w}\|$ is bounded by $\|\boldsymbol{w}\|\leq R$. For a graph $\mathcal{G}(\mathcal{V},\mathcal{E},\boldsymbol{X})\sim ASBM(n_{a},n_{n},\boldsymbol{\mu},\boldsymbol{\nu},(p_{1},q_{1}),(p_{2},q_{2}),P_{a},P_{n})$, when low- and high-pass filters are applied separately to the homophilic and heterophilic node sets $\mathcal{H}_{o},\mathcal{H}_{e}$, there exists an optimal $\boldsymbol{w}^{*},b^{*}$ such that all nodes are linearly separable with the following probability:
$\mathbb{P}\left(\left(\boldsymbol{\hat{X}_{i}}\right)_{i\in\mathcal{V}^{L}}\text{is linearly separable}\right)=1-o_{d}(1),$ (14)
where $o_{d}(1)$ denotes a quantity that converges to 0 as $d\to\infty$.
The proof is provided in Appendix D. Theorem 1 theoretically establishes that, under appropriate conditions, adaptively applying low-pass and high-pass filters to nodes based on local homophily is possible to achieve linear separability across all nodes.
Our architectural design follows this principle by first extracting candidate representations for each node using both low-pass and high-pass filters. These representations are then frozen during the downstream task, aligning with the setting in the theorem where fixed features are used. Subsequently, the gated fusion network adaptively combines the frozen representations. Guided by class-specific fusion preferences and classification loss, our model learns to sense local homophily disparity and adjust fusion weights accordingly. This enables the learned fusion strategy to approximate the node-wise filter assignment in the theorem, allocating representations based on each node’s local homophily, thereby approaching the theoretical bound of linear separability and ultimately improving GAD performance.
@@ 4.1 Experimental Setup
We conduct experiments on 10 GADBench GADBench datasets spanning diverse domains and scales. We compare with a broad range of baseline methods, including standard GNNs GCN ; GIN ; GAT ; ACM ; FAGCN ; AdaGNN ; BernNet , GAD-specific models GAS ; DCI ; PCGNN ; AMNet ; BWGNN ; GHRN ; ConsisGAD ; SpaceGNN , and graph pre-training approaches DGI ; GRACE ; G-BT ; GraphMAE ; BGRL ; SSGE ; PolyGCL . For detailed dataset statistics, dataset descriptions, and baseline descriptions, please kindly see Table 3, Appendix G.1, and Appendix G.2, respectively.
@@
Following GADBench GADBench , we use AUPRC, AUROC, and Rec@K as metrics, with 100 labeled nodes (20 anomalies) per training set. All experiments are averaged over 10 random splits provided by GADBench for robustness. Due to space constraints, detailed evaluation protocols and hyperparameter settings are provided in Appendix G.3 and Appendix G.4, respectively.
Table 1: Comparison of AUPRC for each model. "-" denotes "out of memory". The best and runner-up models are bolded and underlined.
@@ 4.2 Performance Comparison
Overall, graph pre-training methods achieve competitive or even superior performance compared to GAD-specific models, especially on Reddit, Weibo, and Tolokers. In particular, DGI, BWDGI, and APF achieve AUPRC gains of +7.3%, +3.8%, and +4.9% over their end-to-end training counterparts (GCN, BWGNN, and APF (w/o $\mathcal{L}_{pt}$)). These gains highlight the importance of pre-training in addressing the label scarcity problem inherent to GAD, by providing more expressive and transferable initializations.
@@
As observed, APF outperforms both GAD-specific approaches and graph pre-training methods in most cases. On average, APF achieves gains of +6.7% in AUPRC, +3.8% in AUROC, and +6.2% in Rec@K. Even without the pre-training, APF (w/o $\mathcal{L}_{pt}$) surpasses all baseline methods, including BWGNN and AMNet, which also adopt multi-filter architectures. This highlights the strength of our fine-tuning module in adaptively emphasizing anomaly-relevant signals and approximating the theoretical linear separability established in Theorem 1. We note that APF underperforms XGBGraph on Amazon and Elliptic, likely due to the tabular and highly heterogeneous node features that favor tree-based models Tree ; GADBench , as further discussed in Appendix C. Overall, these strong results validate the effectiveness of our proposed anomaly-aware pre-training and granularity-adaptive fine-tuning framework.
@@
To investigate the ability of our model to mitigate the impact of local homophily disparity, we conduct a fine-grained performance analysis on abnormal nodes with varying degrees of local homophily. Specifically, we divide the abnormal nodes in the test set into four quartiles based on their local homophily scores, and compute the model performance within each group against the remaining normal nodes. The results are visualized in Figure 1, 3, 10 and 11.
@@
We observe that detection performance typically declines as local homophily decreases, highlighting the difficulty of identifying anomalies in heterophilic regions and the need for node-adaptive mechanisms. APF consistently achieves stronger performance across all homophily quartiles, demonstrating enhanced robustness to homophily disparity. Such robustness mainly stems from our anomaly-aware pre-training and the adaptive fusion mechanism, allowing the model to tailor its decision boundary to each node’s local structural context and mitigating performance degradation under local homophily disparity.
4.3 Model Analysis
@@
We conduct ablation experiments to examine the impact of key components in both the pre-training and fine-tuning stages of APF. In the pre-training stage, we test three variants: using only the low-pass filter $g_{L}(\cdot)$, only the high-pass filter $g_{H}(\cdot)$, and removing the Rayleigh Quotient-guided sub-tree $\mathcal{G}^{RQ}$ by applying standard DGI on $g_{L}(\cdot)$ and $g_{H}(\cdot)$. In the fine-tuning stage, we evaluate the effect of removing the anomaly-aware regularization loss $\mathcal{L}_{reg}$ and replacing our node- and dimension-adaptive fusion with simpler alternatives such as mean, concatenation, and attention-based fusion AMNet . The results, shown in Table 2, reveal several important findings. (1) Using information from both low- and high-pass filters generally yields better performance than relying on either alone, confirming the necessity of dual-filter pre-training to capture both semantic regularities and anomaly-relevant irregularities in graph structure. (2) Our node- and dimension-adaptive fusion strategy consistently outperforms the alternatives, demonstrating its effectiveness in adapting to node-specific structural variation. (3) The Rayleigh Quotient-guided sub-tree $\mathcal{G}^{RQ}$ and the regularization term $\mathcal{L}_{reg}$ both contribute to improved anomaly discrimination, further boosting performance by encouraging more anomaly-aware representation learning.
@@
Variants YelpChi Questions DGraph-Fin
$g_{L}$ $g_{H}$ $\mathcal{G}^{RQ}$ $\mathcal{L}_{reg}$ Fusion AUPRC AUROC Rec@K AUPRC AUROC Rec@K AUPRC AUROC Rec@K
NDApt 28.4±1.4 68.2±2.3 31.4±1.6 12.3±1.6 71.9±2.1 16.5±0.9 2.9±0.2 72.4±1.3 4.2±0.7
NDApt 27.1±1.6 67.6±1.4 30.2±1.6 10.9±1.8 71.2±2.5 16.0±1.4 2.6±0.1 71.1±0.5 3.3±0.5
- 18.7±1.2 56.5±1.3 20.6±1.7 11.8±1.2 70.6±1.2 16.1±0.9 2.5±0.1 71.0±1.0 3.2±0.1
- 24.4±2.4 64.0±2.5 27.2±2.5 6.3±0.7 66.1±3.2 9.8±1.9 2.6±0.1 71.1±0.5 3.1±0.5
NDApt 27.5±1.6 67.4±2.3 30.6±1.9 11.1±1.6 71.2±2.7 15.9±0.9 2.8±0.1 71.7±0.4 3.7±0.2
Mean 27.6±1.6 67.2±2.1 30.5±1.8 10.8±1.7 71.0±2.4 15.7±1.2 2.8±0.1 71.6±0.4 3.8±0.2
Concat 27.3±1.9 67.3±2.2 30.4±2.1 11.1±1.5 68.5±3.2 15.2±1.7 2.8±0.0 71.5±0.2 3.6±0.3
Atten. 26.2±1.6 66.7±1.7 29.5±1.7 8.3±2.2 67.5±3.5 13.6±2.6 2.8±0.0 71.7±0.3 3.8±0.2
@@
Hyperparameter Analysis.
In addition to the adaptive fusion mechanism, our APF further introduces two hyperparameters, $p_{a}$ and $p_{n}$, which represent the expected preference for low-pass representations in abnormal and normal nodes, respectively. To assess their impact on our performance, we vary these hyperparameters from $0.0$ to $1.0$ in increments of $0.1$. The results are presented in Figure 5, 12 and 13. It is observed that the right half of the heatmap, corresponding to relatively larger $p_{n}$ values, generally outperforms the left half. This aligns with the understanding that normal nodes benefit more from low-pass representations for generic knowledge, due to their strong structural consistency with neighbors. Additionally, the optimal combination of $(p_{a},p_{n})$ always appears in the lower-right half of the heatmap, where $p_{a}\leq p_{n}$. This indicates that abnormal nodes are assigned lower $p_{a}$ values, thus placing greater emphasis on anomaly-indicative components, which better capture their deviation from the local context. These observations are consistent with the intuition behind our adaptive fusion design: normal and abnormal nodes require different emphases of knowledge to maximize discriminability. Overall, $p_{a}$ and $p_{n}$ provide a simple yet effective way that consistently guides APF toward strong and stable anomaly detection performance with minimal tuning effort.
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
Additional Analysis.
Due to space constraints, the additional results, including (1) efficiency analyses regarding training time and memory overhead, (2) performance comparisons under varying numbers of labeled nodes, and (3) additional figurative and tabular results for the presented analyses, are deferred in Appendix F and G.5.

### v3

@@ 3.3 Theoretical Insights
In this subsection, we provide a theoretical analysis to support our architectural design. Our theoretical analysis is grounded in the Contextual Stochastic Block Model (Deshpande et al., 2018), a widely used generative model for attributed graphs (Baranwal et al., 2021; Ma et al., 2022; Mao et al., 2023; Han et al., 2024). To properly reflect homophily disparity, degree heterogeneity, and class imbalance in GAD, we first introduce a variant, the Anomalous Stochastic Block Model (ASBM).
Definition 1 ($ASBM(n_{a},n_{n},\boldsymbol{\mu},\boldsymbol{\nu},(p_{1},q_{1}),(p_{2},q_{2}),\boldsymbol{\theta})$).
Let $\mathcal{C}_{a}$ and $\mathcal{C}_{n}$ be the abnormal and normal node sets with sizes $n_{a}$ and $n_{n}$, respectively; $n=n_{a}+n_{n}$, class priors $\pi_{a}=n_{a}/n$, $\pi_{n}=1-\pi_{a}$ with $\pi_{a}\ll\pi_{n}$. Node features are sampled row-wise as $\boldsymbol{X}_{a}\sim\mathcal{N}(\boldsymbol{\mu},\tfrac{1}{d}\boldsymbol{I})$ and $\boldsymbol{X}_{n}\sim\mathcal{N}(\boldsymbol{\nu},\tfrac{1}{d}\boldsymbol{I})$ with $\|\boldsymbol{\mu}\|_{2},\|\boldsymbol{\nu}\|_{2}\leq 1$, and the full matrix $\boldsymbol{X}=[\boldsymbol{X}_{a};\boldsymbol{X}_{n}]$. Each node $v_{i}$ has a random variable $\theta_{i}>0$ (collected in $\boldsymbol{\theta}$) as the degree parameter, which controls how many edges it tends to form. Nodes can follow either a homophilic or heterophilic connectivity pattern: a node in the homophilic set $\mathcal{H}_{o}$ prefers to connect to nodes of the same class, while a node in the heterophilic set $\mathcal{H}_{e}$ prefers the opposite. Accordingly, edges are generated independently as follows:
[…]
Here $y_{i}\in\{a,n\}$ is the class label and $h_{i}\in\{ho,he\}$ indicates the local pattern (homophilic/heterophilic) adopted by node $v_{i}$; $p_{1}>q_{1}$ enforces homophily (intra-class edges more likely), while $p_{2}<q_{2}$ enforces heterophily (inter-class edges more likely).
In line with prior analyses (Baranwal et al., 2021), we adopt a linear classifier parameterized by $\boldsymbol{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, with predictions $\boldsymbol{\hat{y}}=\sigma(\boldsymbol{\hat{X}}\boldsymbol{w}+b\boldsymbol{1})$ on frozen filtered features $\boldsymbol{\hat{X}}=\boldsymbol{U}g(\boldsymbol{\Lambda})\boldsymbol{U}^{\top}\boldsymbol{X}$, optimized by the binary cross-entropy in Eq. 12. The separability of this model under node-adaptive filtering is characterized by the following theorem:
@@
For a graph $\mathcal{G}(\mathcal{V},\mathcal{E},\boldsymbol{X})\sim ASBM(n_{a},n_{n},\boldsymbol{\mu},\boldsymbol{\nu},(p_{1},q_{1}),(p_{2},q_{2}),\boldsymbol{\theta})$, when low- and high-pass filters are applied separately to the homophilic and heterophilic node sets $\mathcal{H}_{o},\mathcal{H}_{e}$, there exist parameters $\boldsymbol{w}^{*},b^{*}$ such that all nodes are linearly separable with probability $1-o_{d}(1)$.
We present the detailed proof in Appendix E. This theorem theoretically establishes that, under appropriate conditions, adaptively applying low-pass and high-pass filters to nodes based on local homophily is possible to achieve linear separability across all nodes.
While the theorem assumes an idealized scenario with oracle filter assignments, our architectural design implements this principle in a data-driven manner. APF first extracts candidate representations via both low-pass and high-pass filters. Then, instead of hard filter selection, it employs a Gated Fusion Network to generate soft, continuous coefficients that approximate the node-wise filter assignment. Guided by class-specific fusion preferences and classification loss, our model learns to sense local homophily disparity and adjust fusion weights accordingly. This enables the learned fusion strategy to mimic the theoretical node-wise filter assignment without requiring all nodes’ pattern labels, allocating representations based on each node’s local homophily, thereby approaching the theoretical bound of linear separability and ultimately improving GAD performance.
@@ 4 Experiments
In this section, we first evaluate the effectiveness of pre-training for GAD, the overall performance of our model, and its ability to mitigate homophily disparity (Section 4.2). We then analyze the contribution of each component and provide visualizations of the learned fusion coefficients (Section 4.3). Due to space constraints, additional experimental results, such as efficiency comparison, hyperparameter analysis, and representation visualization, are presented in Appendix I.
@@ 4.1 Experimental Setup
We conduct experiments on 10 GADBench (Tang et al., 2023) datasets spanning diverse domains and scales. We compare with a broad range of baseline methods, including standard GNNs (Kipf and Welling, 2017; Xu et al., 2019; Veličković et al., 2018; Luan et al., 2022; Bo et al., 2021; Dong et al., 2021; He et al., 2021), GAD-specific models (Li et al., 2019; Wang et al., 2021; Liu et al., 2021a; Chai et al., 2022; Tang et al., 2022; Gao et al., 2023; Chen et al., 2024b; Dong et al., 2025), and graph pre-training approaches (Veličković et al., 2019; Zhu et al., 2020; Bielak et al., 2022; Hou et al., 2022; Thakoor et al., 2022; Liu et al., 2024c; Chen et al., 2024a). For detailed dataset statistics, dataset descriptions, and baseline descriptions, please kindly see Table 4, Appendix H.1, and Appendix H.2, respectively.
@@
To strictly align with real-world label scarcity, we follow the semi-supervised setting defined in GADBench (Tang et al., 2023). Specifically, we standardize the training set to include exactly 100 labeled nodes (20 anomalies and 80 normal nodes) across all datasets. Thus, the baseline results reported in this paper correspond to the semi-supervised performance metrics found in the GADBench Appendix (Table 11). For evaluation, we employ AUPRC, AUROC, and Rec@K. We prioritize AUPRC over the threshold-dependent F1 score to ensure a robust assessment of precision-recall tradeoffs independent of decision thresholds, which can be unstable under limited supervision. All experiments are averaged over 10 random splits provided by GADBench for robustness. Due to space constraints, detailed evaluation protocols and hyperparameter settings are provided in Appendix H.3 and Appendix H.4, respectively. Our code is available at https://github.com/Cloudy1225/APF.
Table 1: Comparison of AUPRC for each model. ”-” denotes ”out of memory”. The best and runner-up models are bolded and underlined.
@@ 4.2 Performance Comparison
Overall, pre-training methods (from DGI to BWDGI and APF) achieve competitive or even superior performance compared to GAD-specific models, especially on Reddit, Weibo, and Tolokers. In particular, DGI, BWDGI, and APF bring AUPRC improvements of +7.3%, +3.8%, and +4.9% over their respective end-to-end training counterparts (GCN, BWGNN, and APF (w/o $\mathcal{L}_{pt}$)). These gains highlight the importance of pre-training in addressing the label scarcity inherent to GAD, by providing more expressive and transferable initializations.
To further validate the benefits of our pre-training, we also evaluate the learned representations in a purely unsupervised setting, where they are directly coupled with an anomaly scorer IF (Liu et al., 2008) without fine-tuning. The results, presented in Appendix I.1, confirm that our pre-training further enhances unsupervised detection performance, demonstrating its general effectiveness.
@@
As observed, APF outperforms both GAD-specific approaches and graph pre-training methods in most cases. On average, APF achieves gains of +6.7% in AUPRC, +3.8% in AUROC, and +6.2% in Rec@K. Even without the pre-training, APF (w/o $\mathcal{L}_{pt}$) surpasses all baseline methods, including BWGNN and AMNet, which also adopt multi-filter architectures. This highlights the strength of our fine-tuning module in adaptively emphasizing anomaly-relevant signals and approximating the theoretical linear separability established in Theorem 1. We note that APF underperforms XGBGraph on Amazon and Elliptic, likely due to the tabular and highly heterogeneous node features that favor tree-based models (Grinsztajn et al., 2022; Tang et al., 2023), as further discussed in Appendix D. Overall, these strong results validate the effectiveness of our proposed anomaly-aware pre-training and granularity-adaptive fine-tuning framework.
@@
To investigate the ability of our model to mitigate the impact of local homophily disparity, we conduct a fine-grained performance analysis on abnormal nodes with varying degrees of local homophily. Specifically, we divide the abnormal nodes in the test set into four quartiles based on their local homophily scores and compute the model performance
@@
within each group against the remaining normal nodes. The results are visualized in Figure 1, 3, 13 and 14. We observe that detection performance typically declines as local homophily decreases, highlighting the difficulty of identifying anomalies in heterophilic regions and the need for node-adaptive mechanisms. APF consistently achieves stronger performance across all homophily quartiles, demonstrating enhanced robustness to homophily disparity. Such mitigation mainly stems from our anomaly-aware pre-training and the adaptive fusion mechanism, allowing the model to tailor its decision boundary to each node’s local structural context and mitigating performance degradation under local homophily disparity.
To provide a more fine-grained view, we further measure the performance variance across homophily levels by reporting the AUPRC differences between lower-homophily groups (Q2, Q3, Q4) and the highest-homophily group (Q1). The results in Table 2 show that APF generally yields the smallest or near-smallest performance variance across quartiles. This indicates that, although performance degradation remains as local homophily decreases, APF alleviates the disparity more effectively and yields more stable performance under challenging heterophilic conditions.
Table 2: AUPRC differences (Q1 - Q*) across local homophily quartiles. Smaller absolute values indicate lower performance gaps under homophily disparity.
Dataset Weibo T-Finance YelpChi Amazon
Q1-Q2 Q1-Q3 Q1-Q4 Q1-Q2 Q1-Q3 Q1-Q4 Q1-Q2 Q1-Q3 Q1-Q4 Q1-Q2 Q1-Q3 Q1-Q4
AMNet 9.40 32.88 27.32 3.44 30.27 58.66 -0.71 -0.19 0.54 5.42 20.65 47.10
BWGNN 10.73 39.18 44.89 9.47 56.56 86.60 -1.09 -0.94 0.03 8.89 26.14 56.74
ConsisGAD 13.43 43.95 54.96 3.75 25.04 89.25 -1.75 -0.19 -0.06 -3.24 14.63 39.47
APF -1.18 5.72 31.20 2.06 23.36 76.12 -0.50 0.16 0.78 4.18 17.98 45.46
4.3 Model Analyses
@@
We perform a comprehensive ablation study to disentangle the contributions of the major components in both the pre-training and fine-tuning stages of APF. In the pre-training stage, we design four controlled variants: (i) using only the low-pass filter $g_{L}(\cdot)$, (ii) using only the high-pass filter $g_{H}(\cdot)$, (iii) removing the Rayleigh Quotient-guided subgraph $\mathcal{G}^{RQ}$ by applying standard DGI on $g_{L}(\cdot)$ and $g_{H}(\cdot)$, and (iv) replacing $\mathcal{G}^{RQ}$ with a full $k$-hop subgraph “✢”. In the fine-tuning stage, we examine two factors: (i) discarding the anomaly-aware regularization loss $\mathcal{L}_{reg}$, and (ii) replacing our node- and dimension-adaptive fusion with mean pooling, concatenation, or attention-based fusion (Chai et al., 2022).
The results, shown in Table 3, reveal several important findings. (1) Dual-filter necessity. Leveraging both low- and high-pass filters clearly outperforms using either alone, confirming that semantic regularities and anomaly-indicative irregularities must be jointly captured for GAD. (2) Power of $\mathcal{G}^{RQ}$. The Rayleigh Quotient-guided subgraph yields more precise anomaly discrimination than the full $k$-hop variant “✢”, highlighting the Rayleigh Quotient as an effective label-free anomaly indicator. (3) Adaptive fusion advantage. Our node- and dimension-adaptive fusion consistently surpasses the alternatives, showing the importance of exploiting frequency-selective signals while adapting to node-specific contexts. (4) Regularization gains. The anomaly-aware regularization loss $\mathcal{L}_{reg}$ provides additional boosts, aligning representations with class-level fusion bias and further strengthening anomaly detection.
@@
Variants Reddit YelpChi Questions DGraph-Fin
$g_{L}$ $g_{H}$ $\mathcal{G}^{RQ}$ $\mathcal{L}_{reg}$ Fusion AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
NDApt 5.9±0.9 66.8±3.9 28.4±1.4 68.2±2.3 12.3±1.6 71.9±2.1 2.9±0.2 72.4±1.3
NDApt 5.3±0.6 66.1±2.0 27.1±1.6 67.6±1.4 10.9±1.8 71.2±2.5 2.6±0.1 71.1±0.5
NDApt 5.5±0.7 66.5±3.8 27.4±2.8 67.3±2.3 10.8±1.6 71.0±2.3 2.7±0.2 71.0±2.2
- 4.9±0.5 65.3±1.7 18.7±1.2 56.5±1.3 11.8±1.2 70.6±1.2 2.5±0.1 71.0±1.0
- 4.5±0.5 60.1±3.0 24.4±2.4 64.0±2.5 6.3±0.7 66.1±3.2 2.6±0.1 71.1±0.5
NDApt 5.3±0.5 64.3±1.7 27.5±1.6 67.4±2.3 11.1±1.6 71.2±2.7 2.8±0.1 71.7±0.4
Mean 5.1±0.6 63.5±2.0 27.6±1.6 67.2±2.1 10.8±1.7 71.0±2.4 2.8±0.1 71.6±0.4
Concat 5.1±0.6 64.6±2.7 27.3±1.9 67.3±2.2 11.1±1.5 68.5±3.2 2.8±0.0 71.5±0.2
Atten. 5.2±0.7 62.8±2.3 26.2±1.6 66.7±1.7 8.3±2.2 67.5±3.5 2.8±0.0 71.7±0.3
@@
Additional Analyses.
Due to space limitations, please kindly see the Appendix for extended discussions. Specifically, we report the model’s time complexity in Appendix G, and present efficiency analyses regarding training time and memory overhead in Appendix I.3. The effectiveness of our pre-training under a purely unsupervised GAD setting is discussed in Appendix I.1. Moreover, Appendix I.2 provides visualizations of the learned low-pass, high-pass, and fused node representations. We also analyze the sensitivity to hyperparameters $p_{a}$ and $p_{n}$ in Appendix I.4, investigate performance under varying amounts of labeled nodes in Appendix I.5, and include additional figures and tables complementing the main results. We further re-classify the baselines into supervised, semi-supervised, and self-supervised categories in Appendix I.8, and compare our two-stage (pre-training and fine-tuning) approach against a joint learning strategy in Appendix I.9. These analyses jointly provide a more comprehensive understanding of the proposed framework.